# GRAMMAR: Grounded and Modular Methodology for Assessment of Closed-Domain Retrieval-Augmented Language Models

### v6

@@ 4 GRAMMAR
This section introduces GRAMMAR-Gen and GRAMMAR-Eval for grounded and modular evaluation: 1) GRAMMAR-Gen begins by generating templates based on a database schema, enabling scalable data creation  [^3] Further details can be found in Appendices D.2 and E. By leveraging LLMs, GRAMMAR-Gen offers controllability in generating text variations, which not only supports scalable data generation but also facilitates modular robustness evaluation and hypothesis testing. The process of SQL and text template generation is detailed in § 4.1. SQL queries, serving as intermediate representations, are used to create ground-truth answers by querying relational databases, as explained in § 4.2. The overall process is illustrated in Algorithm 1 and Figure 2. 2) GRAMMAR-Eval, an evaluation protocol that effectively pinpoints defective modules by assessing the knowledge deficit of the retrieval database and detecting robustness issues originating from either the retrieval mechanisms or the language models. Further details can be found in § 4.3.
@@ 4.3 Modular Evaluation
The process involves grouping and tagging queries based on their semantics. Specifically, with GRAMMAR-Gen, textual queries $Q$ generated from a particular SQL query $q_{\mathrm{sql}}$ are organized into a semantic group. Depending on the model $M$’s performance, $q_{\mathrm{sql}}$ (and consequently, the group $Q$) are tagged into three principal categories, as shown in Figure 1.
  • Gap Groups emerge when the model $M$ consistently provides incorrect answers for every text query within a subset $S\subseteq Q$ produced through the data generation process. This reflects a deficiency in $M$ ’s knowledge or capability. Formally, this scenario is described as:
@@
    where $M_{e}$ is the evaluation model, which returns “TRUE” if the model’s response $M\left(q_{t}\right)$ is correct.
  • Robust Groups are established when model $M$ accurately responds to all text queries within the subset $S$, evidencing a robust comprehension of the SQL logic. Formally, this is articulated as:
@@
  • Non-robust Groups are characterized by model $M$’s ability to potentially answer the query semantics correctl, but with at least one query $q_{t}$ within $S$ that $M$ fails to predict correctly. This is formally denoted as:
@@
To ensure accurate evaluation of the retrieval and language models without interference from retrieval database errors (i.e., distinguishing knowledge gaps from robustness performance), two strategies have been developed:
  1. 1.
    Removing Gap Examples: To assess the other two modules, we calculate refined accuracy, denoted as $R$, by excluding gap examples:
@@
    where “#” represents “the number of instances”.
  2. 2.
    Balancing Gap Groups: To understand the advantage of this refined accuracy over baseline accuracy Acc, consider the following relationship:
@@ 5.1 Experiment Setup
Closed-domain RAG systems lack established benchmarks, especially because questions must be answered using documents containing knowledge that is not publicly available or known to public LLMs. To address this gap, we modify the Spider benchmark (Yu et al., 2018) and create two synthetic benchmarks with fictitious data. 1) Context-Augmented Spider (Open-Domain): We use the “company_employee” relational database from Spider to simulate a general real-world scenario. Documents for retrieval are synthesized using GPT-4 based on the database. 2) Context-Augmented Spider (Closed-Domain): This is similar to the first dataset, but with company names modified to fictitious entities that are either unused by specific companies or generic according to the latest GPT-4 models. 3) Aurp: To create a strictly closed-domain evaluation, we synthesize fictitious facts about a fictional company called Aurp. A relational database and corresponding retrieval documents are generated so that questions derived from the database can only be answered by the provided documents. Details on the data synthesis process are provided in Appendix G, and the database schemas are outlined in Appendix F. Using these three settings, datasets are generated by GRAMMAR-Gen to validate the design of GRAMMAR-Eval. The statistics of the generated datasets are summarized in Table 2.
@@ 6 Conclusions
Our open-source Python package enables easy reproduction of all results, including the regeneration of experimental data and outcomes from our validation study. Detailed instructions on using the package are provided in Appendix H. Due to data privacy policies, the 198 examples and the RAG system from the sponsoring company are not included.
@@ Appendix E Breakdown of Scalable Data Generation
Appendix F Database Schema Overview
@@
Appendix G Synthetic Processes for Data Generation
G.1 Aurp Setup
@@
G.2 Context-Augmented Spider
@@
Evaluation of Factuality in Generated Documents
To evaluate the factuality of generated documents, a Machine Reading Comprehension (MRC) task is setup, where the generated document is given as context to answer the questions encoding the facts it generate on. Note that the correct answer can only source from the generated document.
Appendix H Reproducing Data and Results
H.1 Reproducing Data
@@
H.2 Reproducing Experiment Results

### v7

@@ 4 GRAMMAR
This section introduces GRAMMAR-Gen and GRAMMAR-Eval for grounded and modular evaluation: 1) GRAMMAR-Gen begins by generating templates based on a database schema, enabling scalable data creation  [^3] Further details can be found in Appendices D.2 and E. By leveraging LLMs, GRAMMAR-Gen offers controllability in generating text variations, which not only supports scalable data generation but also facilitates modular robustness evaluation and hypothesis testing. The process of SQL and text template generation is detailed in § 4.1. SQL queries, serving as intermediate representations, are used to create ground-truth answers by querying relational databases, as explained in § 4.2. The overall process is illustrated in Algorithm 1 and Figure 2. Note that the use of relational databases does not limit GRAMMAR’s applicability in practice, as specified in Appendix F. 2) GRAMMAR-Eval, an evaluation protocol that effectively pinpoints defective modules by assessing the knowledge deficit of the retrieval database and detecting robustness issues originating from either the retrieval mechanisms or the language models. Further details can be found in § 4.3.
@@ 4.3 Modular Evaluation
The process involves grouping and tagging queries based on their semantics. Specifically, with GRAMMAR-Gen, textual queries $Q$ generated from a particular SQL query $q_{\mathrm{sql}}$ are organized into a semantic group. Depending on the model $M$’s performance, $q_{\mathrm{sql}}$ (and consequently, the group $Q$) are tagged into three principal categories, as shown in Figure 1: 1) Gap Groups emerge when the model $M$ consistently provides incorrect answers for every text query within a subset $S\subseteq Q$ produced through the data generation process. This reflects a deficiency in $M$ ’s knowledge or capability. Formally, this scenario is described as:
@@
where $M_{e}$ is the evaluation model, which returns “TRUE” if the model’s response $M\left(q_{t}\right)$ is correct. 2) Robust Groups are established when model $M$ accurately responds to all text queries within the subset $S$, evidencing a robust comprehension of the SQL logic. Formally, this is articulated as:
@@
3) Non-robust Groups are characterized by model $M$’s ability to potentially answer the query semantics correctl, but with at least one query $q_{t}$ within $S$ that $M$ fails to predict correctly. This is formally denoted as:
@@
To ensure accurate evaluation of the retrieval and language models without interference from retrieval database errors (i.e., distinguishing knowledge gaps from robustness performance), two strategies have been developed: 1) Removing Gap Examples: To assess the other two modules, we calculate refined accuracy, denoted as $R$, by excluding gap examples:
@@
where “#” represents “the number of instances”. 2) Balancing Gap Groups: To understand the advantage of this refined accuracy over baseline accuracy Acc, consider the following relationship:
@@ 5.1 Experiment Setup
Closed-domain RAG systems lack established benchmarks, especially because questions must be answered using documents containing knowledge that is not publicly available or known to public LLMs. To address this gap, we modify the Spider benchmark (Yu et al., 2018) and create two synthetic benchmarks with fictitious data. 1) Context-Augmented Spider (Open-Domain): We use the “company_employee” relational database from Spider to simulate a general real-world scenario. Documents for retrieval are synthesized using GPT-4 based on the database. 2) Context-Augmented Spider (Closed-Domain): This is similar to the first dataset, but with company names modified to fictitious entities that are either unused by specific companies or generic according to the latest GPT-4 models. 3) Aurp: To create a strictly closed-domain evaluation, we synthesize fictitious facts about a fictional company called Aurp. A relational database and corresponding retrieval documents are generated so that questions derived from the database can only be answered by the provided documents. Details on the data synthesis process are provided in Appendix H, and the database schemas are outlined in Appendix G. Using these three settings, datasets are generated by GRAMMAR-Gen to validate the design of GRAMMAR-Eval. The statistics of the generated datasets are summarized in Table 2. We have evaluated various perspectives on data quality to ensure the reliability of our framework, as detailed in Appendix I.
@@ 6 Conclusions
Our open-source Python package enables easy reproduction of all results, including the regeneration of experimental data and outcomes from our validation study. Detailed instructions on using the package are provided in Appendix J. Due to data privacy policies, the 198 examples and the RAG system from the sponsoring company are not included.
@@ Appendix E Breakdown of Scalable Data Generation
Appendix F Use of Relational Databases Does Not Limit GRAMMAR’s Applicability in Practice
The key point is that SQL databases can be easily self-constructed with a limited understanding of the RAG system’s use cases to construct a simple schema (e.g., Figure 3) and insert a small number of rows into the tables (e.g., only 49 rows are manually inserted across 3 Aurp tables). Our approach is highly scalable for generating evaluation data, as shown in Table 2 and Figure 2, and highlighted in Line 241 with further details in Appendix E.
The requirements of defining a basic schema or inserting rows can be easily handled either by developers who normally have concrete ideas during RAG system development, by probing data during short-term testing of the RAG system before deployment, or even through a brief (1-hour) interview with users.
Appendix G Database Schema Overview
@@
Appendix H Synthetic Processes for Data Generation
H.1 Aurp Setup
@@
H.2 Context-Augmented Spider
@@
Appendix I Evaluation of Data Quality
Evaluation Perspectives
To make our evaluation legitimate in terms of data quality, the following conditions should be satisfied:
  • Perspective 1: The generated text queries ask about the correct fact
  • Perspective 2: The generated documents (for retrieval) contain those facts
  • Perspective 3: Coherence and grammatical correctness are important aspects of data quality. However, it is widely acknowledged that GPT-4 models consistently generate coherent text (Coyne et al., 2023), and our random manual checks further confirm this.
No Human Preferences in Factual QA Data
Our setting is focused on factual QA tasks, where answers are either correct or incorrect, leaving no room for subjective human preferences.
I.1 Evaluating Perspective 1, 2
To verify these two points, a Machine Reading Comprehension (MRC) task is setup, where the generated document is given as context to answer the questions encoding the facts it generate on. In this setup, the queries can only be answered correctly if 1)the given MRC context (i.e., generated documents) contains correct facts, and 2) the queries do ask about the facts (since the MRC answers tend to be incorrect given the large amount of information in the documents).
These two points correspond to verifying condition 2 and 1, respectively. In simple terms, if the MRC model (or humans) consistently generates correct answers, the data quality satisfies conditions 1 and 2.
Empirical Results
Using the balanced, short-query version of Spider (closed-domain) and Aurp datasets, where correct documents are given as context, the MRC model based on GPT-3.5 generates correct answers 80.11% and 98.52% of the time, respectively. This only confirms the data quality to some extent.
Further Analysis
To ensure a full reliability, we did error checking on the remaining 19.89% and 1.48% examples and check whether they undermine GRAMMAR’s reliability. The results show that: All incorrect predictions resulted from failing condition 2: the lack of expected facts in the retrieval documents. However, this does not undermine GRAMMAR’s reliability, since these examples were all correctly identified as gap examples.
Appendix J Reproducing Data and Results
J.1 Reproducing Data
@@ I.1 Evaluating Perspective 1, 2
J.2 Reproducing Experiment Results